# Interpretable and Equation-Free Response Theory for Complex Systems

### v1

@@ II.1 Linear Response
Let’s define $\langle\Psi,\mu\rangle=\sum_{i=1}^{N}\Psi_{i}\mu_{i}$ the the expectation value of an observable $\Psi\in\mathbb{R}^{1\times N}$ according to the measure $\mu$. We then have $\langle\Psi,\nu(n)\rangle=\langle\Psi,{\nu_{inv}}\rangle+\epsilon\langle\Psi,\nu^{(1)}_{i}(n)\rangle+h.o.t.$. We then have:
@@
where we have broken down the Green’s function into the sum of $N-1$ terms, each associated with a specific mode of variability of the system. Note that the first term $i=1$ in the summation given in Eq. 7 vanishes because it can be proved that $\Pi^{T}_{1}m=0$, see Lucarini (2016). Since $\lambda_{i}^{k}=\exp(k\nu_{i})$, with $\Re[\nu_{i}]<0$, the previous expansion provides a specific statistical model - the sum of exponentials - for fitting a Green’s function from data. We will comment on this matter in Sect. V. The results presented in Eqs. 7-8 correspond, in the case of a finite-state Markov chain, to the key findings shown in Santos Gutiérrez and Lucarini (2022) for a general Langevin equation. The derivation is much simpler in the case presented here whilst very little is lost at conceptual level. If one assumes that $f=1$ and takes the l
@@ II.2 Second order response
Note again that our summation excludes the term corresponding to the invariant measure of the systen. The results above can easily be extended at all orders of perturbations, see Appendix A. This implies that the full nonlinear time-dependent response can be obtained from the knowledge of $m$, $\mathcal{M}$, and $f$ for any observable of the system. Assuming $f=1$ and taking the $n\rightarrow\infty$ limit, the results presented in Lucarini (2016) are easily recovered.
@@ III Response theory for Correlations
In the vast majority of cases, response theory has been used to computate the change of the measure of a system resulting from applied forcings. Yet, in many practical cases, it is relevant to study the impact of the perturbation on the temporal correlation properties of the system. Taking the example of climate science to illustrate this point, response theory applied to observables describes the change in the state of the climate at a certain time horizon with respect to a reference climatology, whilst response theory applied to correlation describes how climate variability and so-called teleconnections like the North Atlantic Oscillation or El-Niño-Southern Oscillation are impacted by the applied perturbation Ghil and Lucarini (2020); Lucarini and Chekroun (2023b). Some preliminary contributions to the development of response formulas for correlations have been presented in Lucarini and Wouters (2017), which provides a reference for the results presented below.
Let us define $C_{l}(\Psi,\Phi)=\langle(\mathcal{M}^{T})^{l}\Psi\circ\Phi,\nu_{inv}\rangle-\langle\Psi,\nu_{inv}\rangle\langle\Phi,\nu_{inv}\rangle$ the unperturbed $l-$lagged correlation between the two observables $\Psi$ and $\Phi$, where $\circ$ indicates the Hadamard product. We first consider the case of a static perturbation to the Markov process of the form $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon}=\mathcal{M}+\epsilon m$, which has not yet been explicitly treated in the literature up to our knowledge. We then have $C^{\epsilon}_{l}(\Psi,\Phi)=\langle(\mathcal{M}^{T}+\epsilon m^{T})^{l}\Psi\Phi,\nu_{\epsilon}\rangle-\langle\Psi,\nu_{\epsilon}\rangle\langle\Phi,\nu_{\epsilon}\rangle$. Now we write down the terms up to first order in $\epsilon$ in the expression of the correlation. For the first term, we obtain:
@@
Hence, the sensitivity of a time lagged correlation between $\Psi$ and $\Phi$ is nontrivial and can be broken up into four terms. The first term $\delta^{(1)}_{a,\epsilon}(\Psi(l),\Phi)$ is associated with the impact of perturbation on the evolution law of the system in the time interval of length $l$. The second term $\delta^{(1)}_{b,\epsilon}(\Psi(l),\Phi)$ describes the change in the expectation value of the product $\Psi(l),\Phi$ due to the variation of the invariant measure. The last two terms associated with the change in the expectation value of the two observables. Note that the matrix expression provided here is much simpler than the corresponding functional expression given in Lucarini and Wouters (2017). We remark that by performing the spectral expansion of the $\mathcal{M}^{T}$ matrix we are able to disentangle the contributions coming from the various modes of the Koopman operator.
It is indeed possible to extend the response theory for correlation to the case where $f$ has a non-trivial time dependence. The results are reported in Appendix B. It is important to note that the resulting formulas allow us to define how correlations behave in a non-autonomous system. In this case, correlations at time $t$ need to be interpreted as integrals performed on the measure of the snapshot attractor at time $t$ (which is a slice of the pullback attractor Ghil et al. (2008); Chekroun et al. (2011)); see discussion in Bódai et al. (2011); Tél et al. (2020); Bódai et al. (2020).
@@ IV A Simple Example
The example is taken from Klus et al. (2016) with a small modification. The potential is depicted in Fig. 1a and features three local minima, located at $(x_{1},y_{1})\approx(0,1.55)$, $(x_{2},y_{2})=(-1,0)$, and $(x_{3},y_{3})=(1,0)$. The potential we consider here corresponds almost exactly to a case study presented in Klus et al. (2016). We have included an additional term $-y$ in the definition of the potential in order to lower the local minimum in $(x_{1},y_{1})$ to a value closer to the absolute minimum that is realized at $(x_{2},y_{2})$ and $(x_{3},y_{3})$.
@@
Figure 1: (a) Potential function $V(x,y)$ with approximate indication of the three quasi-invariant regions surrounding the minina of the $V$.(b) Invariant Measure $\rho_{0}\propto\exp(-2V(x,y)/\sigma^{2})$. (c). First subdominant Kolmogorov mode, $\lambda_{1}=0.9916$. (d) First subdominant mode of the Perron-Frobenius Operator. (e) Second subdominant Kolmogorov model, $\lambda_{2}=0.9655$. (f) Second subdominant mode of the Perron-Frobenius Operator.
@@
We have that $P^{\tau}=\exp(\mathcal{L}\tau)$ is the Perron-Frobenius operator describing the evolution of measures for a time $\tau$. The unique invariant measure in depicted in Fig. 1b, obeys $\mathcal{L}_{0}\rho_{0}=0$, and is given by $\rho_{0}=Z^{-1}\exp(-2V/\sigma^{2})$, where $Z$ is the normalisation factor. In what follows, we assume $\sigma=0.7$. Our numerical simulations are performed using the standard Euler-Maruyama scheme Kloeden and Platen (1992) with $dt=0.05$. We sample our output every 20 times steps, so that our reference time scale is $\tau=1$. The dynamics given by Eqs. 23-26 describes an equilibrium system.
@@
To approximate $P^{\tau}$, we discretize the phase space of the system into $N$ disjoint subsets $\{B_{1},B_{2},\dots,B_{N}\}$, forming a partition of the space. The Ulam transfer operator $P^{\tau}_{\{N\},ij}$, a finite-dimensional stochastic matrix describes the probability that the orbit of the system is at time $t+\tau$ in the subset $B_{i}$ is at time $t$ it was in the subset $B_{j}$. In our case, the $B_{j}$’s are given by the 625 cubes with side 0.2 centred around the origin. Whilst in principle one would need to cover the entire $\mathbb{R}^{2}$ the confining potential makes all regions beyond those we consider here entirely irrelevant unless one consider extremely long time scales.
@@
Following Klus et al. (2016), we populate each cube with 1000 ensemble members distributed uniformly according to the Lebesgue measure. Each member evolves for $\tau=1$ time unit. We then construct an estimate of $P^{\tau}_{\{N\}}$ by counting the transitions. We repeat the operation 20 times and by averaging we obtain our best estimate of $P^{\tau}_{\{N\}}$, which constitutes our reference discretized stochastic matrix, so that $\mathcal{M}=P^{\tau}_{\{N\}}$. The solution to $\mathcal{M}\nu_{inv}=\nu_{inv}$ gives an extremely good approximation in the gridded domain defined by the $B_{j}$’s of the true invariant measure $\rho_{0}$. The discretized system obeys detailed balance, as expected from the nature of the original continuum system. The corresponding left eigenvector, which is the first right eigenvector of the Koopman operator $\mathcal{M}^{T}$, is constant everywhere. We obtain 624 additional eigenvalue-eigenvector pairs for the matrix $\mathcal{M}$. Since the system obeys detailed balance, we would expect all of them to be real, but, because of insufficient sampling and numerical precision, this constraint is not necessarily obeyed for the rapidly dacaying ones.
@@
The two subdominant eigenvectors are depicted in Fig. 1d and Fig. 1f. The first one describes the transitions between the neighbouhoods of $(x_{2},y_{2})$ and $(x_{3},y_{3})$. The second one describes the transitions between the neighbouhoods of $(x_{1},y_{1})$ and either $(x_{2},y_{2})$ or $(x_{3},y_{3})$. The corresponding subdominant eigenvectors of the Koopman operator are depicted in Fig. 1c and Fig. 1e. There is a very large spectral gap between the three dominant modes (corresponding to $\lambda=1$, $\lambda\approx 0.9917$, $\lambda\approx 0.9655$ the rest of the spectral components ($\lambda_{4}\approx 0.050)$, which indicates that the system can safely undergo a model reduction procedure. Indeed, the level sets of the two Koopman modes can be used to identify the three metastable regimes of the system, which are indicated as (1), (2) and (3) in Fig. 1a and correspond closely to the basins of attraction of the three local minima in the case the stochastic forcing is switched off.
@@
We can introduce a more efficient reduced-order representation of the dynamics, whereby the phase space of the system is partitioned into just three states, corresponding to the regions (1), (2), and (3). In the simple case described here, this corresponds to MSM. This amounts to neglecting entirely the intrawell dynamics, which, in the Ulam description above, is mostly captured by the Koopman modes with index $\geq 4$. We estimate the reduced Markov model (RMM) discretized transfer operator $\tilde{M}$ by performing a single run lasting $10^{7}$ time units (after disregarding a small transient) and counting the transitions between the 3 states described above. The obtained estimates for $\tilde{M}$ and its eigenvectors are given below.
$\tilde{M}=\left(\begin{array}[]{ccc}0.9701&0.0095&0.0095\\ 0.0149&0.9904&0.0001\\ 0.0149&0.0001&0.9904\end{array}\right)\quad\nu_{inv}=\begin{pmatrix}0.240\\ 0.380\\ 0.380\end{pmatrix}\quad\nu_{2}=\begin{pmatrix}0.000\\ 0.701\\ -0.701\end{pmatrix}\quad\nu_{3}=\begin{pmatrix}-0.810\\ 0.405\\ 0.405\end{pmatrix}$
@@
Also this minimal model obeys detailed balance and the two nontrivial eigenvalues $\lambda_{2}=0.9904$ and $\lambda_{3}=0.9606$ correspond closely to the first two subdominant eigenvalues obtained using the Ulam method, and the corresponding eigenvectors provide a coarse grained version of the figures provided in Fig. 1d and Fig. 1f.
@@ IV.1 Applying Extra Forcings
are the causal Green’s functions for the observable $\Psi$ associated with the perturbations acting along the $x$ and $y$ directions, respectively. These formulas can be readily derived from the general version of the FDT Marconi et al. (2008); Santos Gutiérrez and Lucarini (2022).
@@ IV.1.1 Linear Response - Observables
In order to estimate the matrices $m_{x}$ for the Ulam discretization, we repeat the same protocol considering the perturbed dynamics realised by choosing $\epsilon_{1}=0.05$ and $f=1$. We derive $P^{\tau,+}_{\{N\}}$. We repeat the experiment by choosing $\epsilon_{1}=-0.05$ and $f=1$, and derive $P^{\tau,-}_{\{N\}}$. We estimate $m_{x}=(P^{\tau,+}_{\{N\}}-P^{\tau,-}_{\{N\}})/(2\epsilon_{1})$. Note that using centred differences ensures high precisions when studying linear response Gritsun and Lucarini (2017). In order to estimate $m_{y}$, we repeat the same procedure described above by considering $\epsilon_{2}=-0.05$ and $\epsilon_{2}=-0.05$, respectively.
@@
The estimates we obtain for the Green’s functions of interest are shown in Fig. 2. When considering RMM, we clearly see that $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$, which as a result of the choice of forcing/observable pair, only one Koopman mode is retained in the expansion given in Eqs. 7- 8. Indeed, it is clear that if we force along the $x$-direction and choose $x$ as observable, the second Koopman mode (or only the second Perron-Frobenium mode) is the only one that matters. Similarly, the third mode is the only one retained in the spectral expansion of the Green’s function when forcing along $y$-direction and choosing $y$ as observable. Remarkably, also when considering the much higher complexity Markov chain constructed through the Ulam method, to a very good approximation we have $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$ (note that the corresponding $\lambda$’s are slightly larger than in the RMM case). This means that by and large only one of the 625 natural modes of variability of the system matters in defining the response to perturbation, at least in the cases we consider here. This shows that the Koopman decomposition provides the much desired property of interpretability of the response.
! Refer to caption
Figure 2: Linear Green’s function for the $x$ and $y$ observables for additive forcing acting on $x$ (index $m_{x}$) or $y$ (index $m_{y}$) direction. Results are shown for the FDT estimate and the estimates obtained using Markov models constructed with Ulam’s discretization and the 3-state RMM. The inset emphasizes teh exponential decay of the Green’s functions.

### v2

@@ II.1 Linear Response
Let us define $\langle\Psi,\mu\rangle=\sum_{i=1}^{N}\Psi_{i}\mu_{i}$ the the expectation value of an observable $\Psi\in\mathbb{R}^{1\times N}$ according to the measure $\mu$. We then have $\langle\Psi,\nu(n)\rangle=\langle\Psi,{\nu_{inv}}\rangle+\epsilon\langle\Psi,\nu^{(1)}_{i}(n)\rangle+h.o.t.$. We then have:
@@
where we have broken down the Green’s function into the sum of $N-1$ terms, each associated with a specific mode of variability of the system. Note that the first term $i=1$ in the summation given in Eq. 7 vanishes because it can be proved that $\Pi^{T}_{1}m=0$, see [41]. Since $\lambda_{i}^{k}=\exp(k\nu_{i})$, with $\Re[\nu_{i}]<0$, the previous expansion provides a specific statistical model - the sum of exponentials - for fitting a Green’s function from data. We will comment on this matter in Sect. V. The results presented in Eqs. 7-8 correspond, in the case of a finite-state Markov chain, to the key findings shown in [17] for a general Langevin equation. The derivation is much simpler in the case presented here whilst very little is lost at conceptual level. If one assumes that $f=1$ and takes the l
@@ II.2 Second order response
Note again that our summation excludes the term corresponding to the invariant measure of the systen. The results above can easily be extended at all orders of perturbations, see Appendix A. This implies that the full nonlinear time-dependent response can be obtained from the knowledge of $m$, $\mathcal{M}$, and $f$ for any observable of the system. Assuming $f=1$ and taking the $n\rightarrow\infty$ limit, the results presented in [41] are easily recovered.
@@ III Response theory for Correlations
In the vast majority of cases, response theory has been used to computate the change of the measure of a system resulting from applied forcings. Yet, in many practical cases, it is relevant to study the impact of the perturbation on the temporal correlation properties of the system. Taking the example of climate science to illustrate this point, response theory applied to observables describes the change in the state of the climate at a certain time horizon with respect to a reference climatology, whilst response theory applied to correlation describes how climate variability and so-called teleconnections like the North Atlantic Oscillation or El-Niño-Southern Oscillation are impacted by the applied perturbation [54, 55]. Some preliminary contributions to the development of response formulas for correlations have been presented in [48], which provides a reference for the results presented below.
Let us define $C_{l}(\Psi,\Phi)=\langle(\mathcal{M}^{T})^{l}\Psi\circ\Phi,\nu_{inv}\rangle-\langle\Psi,\nu_{inv}\rangle\langle\Phi,\nu_{inv}\rangle$ the unperturbed $l-$lagged correlation between the two observables $\Psi$ and $\Phi$, where $\circ$ indicates the Hadamard product. We first consider the case of a static perturbation to the Markov process of the form $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon}=\mathcal{M}+\epsilon m$, which has not yet been explicitly treated in the literature up to our knowledge. We then have $C^{\epsilon}_{l}(\Psi,\Phi)=\langle(\mathcal{M}^{T}+\epsilon m^{T})^{l}\Psi\circ\Phi,\nu_{\epsilon}\rangle-\langle\Psi,\nu_{\epsilon}\rangle\langle\Phi,\nu_{\epsilon}\rangle$. Now we write down the terms up to first order in $\epsilon$ in the expression of the correlation. For the first term, we obtain:
@@
Hence, the sensitivity of a time lagged correlation between $\Psi$ and $\Phi$ is nontrivial and can be broken up into four terms. The first term $\delta^{(1)}_{a,\epsilon}(\Psi(l),\Phi)$ is associated with the impact of perturbation on the evolution law of the system in the time interval of length $l$. The second term $\delta^{(1)}_{b,\epsilon}(\Psi(l),\Phi)$ describes the change in the expectation value of the product $\Psi(l),\Phi$ due to the variation of the invariant measure. The last two terms associated with the change in the expectation value of the two observables. Note that the matrix expression provided here is much simpler than the corresponding functional expression given in [48]. We remark that by performing the spectral expansion of the $\mathcal{M}^{T}$ matrix we are able to disentangle the contributions coming from the various modes of the Koopman operator.
It is indeed possible to extend the response theory for correlation to the case where $f$ has a non-trivial time dependence. The results are reported in Appendix B. It is important to note that the resulting formulas allow us to define how correlations behave in a non-autonomous system. In this case, correlations at time $t$ need to be interpreted as integrals performed on the measure of the snapshot attractor at time $t$ (which is a slice of the pullback attractor [56, 57]); see discussion in [58, 59, 60].
@@ IV A Simple Example
The example is taken from [49] with a small modification. The potential is depicted in Fig. 1a and features three local minima, located at $(x_{1},y_{1})\approx(0,1.55)$, $(x_{2},y_{2})=(-1,0)$, and $(x_{3},y_{3})=(1,0)$. The potential we consider here corresponds almost exactly to a case study presented in [49]. We have included an additional term $-y$ in the definition of the potential in order to lower the local minimum in $(x_{1},y_{1})$ to a value closer to the absolute minimum that is realized at $(x_{2},y_{2})$ and $(x_{3},y_{3})$.
@@
Figure 1: (a) Potential function $V(x,y)$ with approximate indication of the three quasi-invariant regions surrounding the minina of the $V$.(b) Invariant Measure $\rho_{0}\propto\exp(-2V(x,y)/\sigma^{2})$. (c). First subdominant Kolmogorov mode, $\lambda_{1}=0.9916(2)$. (d) First subdominant mode of the Perron-Frobenius Operator. (e) Second subdominant Kolmogorov model, $\lambda_{2}=0.9655(2)$. (f) Second subdominant mode of the Perron-Frobenius Operator.
@@
where $\mathcal{L}_{0}$ is the Fokker-Planck operator. We have that $P^{\tau,0}=\exp(\mathcal{L}_{0}\tau)$ is the associated Perron-Frobenius operator describing the evolution of measures for a time $\tau$. We refer to the The unique invariant measure in depicted in Fig. 1b, obeys $\mathcal{L}_{0}\rho_{0}=0$, and is given by $\rho_{0}=Z^{-1}\exp(-2V/\sigma^{2})$, where $Z$ is the normalization factor. In what follows, we assume $\sigma=0.7$. Our numerical simulations are performed using the standard Euler-Maruyama scheme [61] with $dt=0.05$. We sample our output every 20 times steps, so that our reference time scale is $\tau=1$. The dynamics given by Eqs. 23-26 describes an equilibrium system.
@@
To approximate $P^{\tau,0}$, we discretize the phase space of the system into $N$ disjoint subsets $\{B_{1},B_{2},\dots,B_{N}\}$, forming a partition of the space. The Ulam transfer operator $P^{\tau,0}_{\{N\},ij}$, a finite-dimensional stochastic matrix describes the probability that the orbit of the system is at time $t+\tau$ in the subset $B_{i}$ is at time $t$ it was in the subset $B_{j}$. In our case, the $B_{j}$’s are given by the 625 cubes with side 0.2 centred around the origin. Whilst in principle one would need to cover the entire $\mathbb{R}^{2}$ the confining potential makes all regions beyond those we consider here entirely irrelevant unless one consider extremely long time scales.
@@
Following [49], we populate each cube with 1000 ensemble members distributed uniformly according to the Lebesgue measure. Each member evolves for $\tau=1$ time unit. We then construct an estimate of $P^{\tau,0}_{\{N\}}$ by counting the transitions. We repeat the operation 20 times and by averaging we obtain our best estimate of $P^{\tau,0}_{\{N\}}$, which constitutes our reference discretized stochastic matrix, so that $\mathcal{M}=P^{\tau,0}_{\{N\}}$. The solution to $\mathcal{M}\nu_{inv}=\nu_{inv}$ gives an extremely good approximation in the gridded domain defined by the $B_{j}$’s of the true invariant measure $\rho_{0}$. The discretized system obeys detailed balance, as expected from the nature of the original continuum system. The corresponding left eigenvector, which is the first right eigenvector of the Koopman operator $\mathcal{M}^{T}$, is constant everywhere. We obtain 624 additional eigenvalue-eigenvector pairs for the matrix $\mathcal{M}$. Since the system obeys detailed balance, we would expect all of them to be real, but, because of insufficient sampling and numerical precision, this constraint is not necessarily obeyed for the rapidly decaying ones.
@@
The two subdominant eigenvectors are depicted in Fig. 1d and Fig. 1f. The first one describes the transitions between the neighbourhoods of $(x_{2},y_{2})$ and $(x_{3},y_{3})$. The second one describes the transitions between the neighbourhoods of $(x_{1},y_{1})$ and either $(x_{2},y_{2})$ or $(x_{3},y_{3})$. The corresponding subdominant eigenvectors of the Koopman operator are depicted in Fig. 1c and Fig. 1e. There is a very large spectral gap between the three dominant modes (corresponding to $\lambda=1$, $\lambda\approx 0.9917(2)$, and $\lambda\approx 0.9655(2)$) and the rest of the spectral components ($\lambda_{4}\approx 0.050(2))$, which indicates that the system can safely undergo a model reduction procedure. Note that the uncertainty on the least significant digit has been estimated using 10 separate integrations. The model reduction could be algebraically achieved by substituting $\mathcal{K}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}$, and $\mathcal{M}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}^{T}$. Additionally, the level sets of the two subdominant Koopman modes can be used to identify the three metastable regimes of the system, which are indicated as (1), (2) and (3) in Fig. 1a and correspond to the basins of attraction of the three local minima in the case the stochastic forcing is switched off.
@@
We use such a geometrical partition of the phase space to introduce a separate reduced-order representation of the dynamics, whereby the phase space of the system is partitioned into three states, corresponding to the regions (1), (2), and (3). In the simple case described here, this corresponds to MSM. This amounts to neglecting entirely the intrawell dynamics, which, in the Ulam description above, is mostly captured by the Koopman modes with index $\geq 4$. We estimate the reduced Markov model (RMM) discretized transfer operator $\tilde{M}$ by performing a single run lasting $10^{7}$ time units (after disregarding a small transient) and counting the transitions between the 3 states described above. The obtained estimates for $\tilde{M}$ and its eigenvectors are given below, where all the numbers indicated below have an approximate uncertainty of 2 in the least significant we have written out.
$\tilde{M}=\left(\begin{array}[]{ccc}0.9701&0.0095&0.0095\\ 0.0149&0.9904&0.00006\\ 0.0149&0.00006&0.9904\end{array}\right)\quad\nu_{inv}=\begin{pmatrix}0.240\\ 0.380\\ 0.380\end{pmatrix}\quad\nu_{2}=\begin{pmatrix}0.000\\ 0.701\\ -0.701\end{pmatrix}\quad\nu_{3}=\begin{pmatrix}-0.810\\ 0.405\\ 0.405\end{pmatrix}$
@@
Also this minimal model obeys detailed balance and the two nontrivial eigenvalues $\lambda_{2}=0.9904(2)$ and $\lambda_{3}=0.9606(2)$ correspond closely to the first two subdominant eigenvalues obtained using the Ulam method, and the corresponding eigenvectors provide a coarse grained version of the figures provided in Fig. 1d and Fig. 1f.
@@ IV.1 Applying Extra Forcings
are the causal Green’s functions for the observable $\Psi$ associated with the perturbations acting along the $x$ and $y$ directions, respectively. These formulas can be readily derived from the general version of the FDT [6, 17].
@@ IV.1.1 Linear Response - Observables
In order to estimate the matrices $m_{x}$ for the Ulam discretization, we repeat the same protocol considering the perturbed dynamics realised by choosing $\epsilon_{1}=0.05$ and $f=1$. We derive $P^{\tau,+}_{\{N\}}$. We repeat the experiment by choosing $\epsilon_{1}=-0.05$ and $f=1$, and derive $P^{\tau,-}_{\{N\}}$. We estimate $m_{x}=(P^{\tau,+}_{\{N\}}-P^{\tau,-}_{\{N\}})/(2\epsilon_{1})$. Note that using centred differences ensures high precisions when studying linear response [62]. In order to estimate $m_{y}$, we repeat the same procedure described above by considering $\epsilon_{2}=-0.05$ and $\epsilon_{2}=-0.05$, respectively.
@@
The estimates we obtain for the Green’s functions of interest are shown in Fig. 2. When considering RMM, we clearly see that $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$, which as a result of the choice of forcing/observable pair, only one Koopman mode is retained in the expansion given in Eqs. 7- 8. Indeed, it is clear that if we force along the $x$-direction and choose $x$ as observable, the second Koopman mode (or only the second Perron-Frobenium mode) is the only one that matters. Similarly, the third mode is the only one retained in the spectral expansion of the Green’s function when forcing along $y$-direction and choosing $y$ as observable.
Remarkably, also when considering the much higher complexity Markov chain constructed through the Ulam method, to a very good approximation we have $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$ (note that the corresponding $\lambda$’s are slightly larger than in the RMM case). This means that by and large only one of the 625 natural modes of variability of the system matters in defining the response to perturbation, at least in the cases we consider here. This shows that the Koopman decomposition provides the much desired property of interpretability of the response.
Figure 2: Linear Green’s function for the $x$ and $y$ observables for additive forcing acting on $x$ (index $m_{x}$) or $y$ (index $m_{y}$) direction. Results are shown for the FDT estimate and the estimates obtained using Markov models constructed with Ulam’s discretization and the 3-state RMM. The inset emphasizes the exponential decay of the Green’s functions.